OpenStudy (anonymous):

MEDAL AND FAN Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -1/2x + 5. y = -2x + 2 y = 2x - 14 y = 1/2x - 8 y = -1/2x - 4

OpenStudy (lucaz):

parallel lines have the same slopes, as you can see the slope of the last equation is -1/2, the same of the line y = -1/2 + 5 also, you have the point (4,-6), so in the standard form you have \[\large{ y-y_1=m(x-x_1) \\ y-(-6)=(-\frac{1}{2})(x-4) \\ y+6=-\frac{1}{2}x+\frac{4}{2} \\ y=-\frac{1}{2}x+2-6 \\ y=-\frac{1}{2}x-4 }\]

OpenStudy (anonymous):

Thank you so much ! Can you help me with a couple more ?

OpenStudy (lucaz):

sorry, I have to go right now, somebody else will help you